# SimSUM – Simulated Benchmark with Structured and Unstructured Medical Records

### v3

@@ 2.1 Modeling structured tabular variables with a Bayesian network
The coefficients for each model were tuned using gradient descent based on the train cases shown in Table 8 in Appendix A. Like before, the expert was asked to (loosely) label these cases with how long they suspected the patient to stay home as a result of these symptoms. The coefficient for the variable self-employed was tuned manually, based on the assumption that being self-employed would shave some days off the predicted number, regardless of the particular symptoms experienced by the patient. As a sanity check, we compared the mean number of days predicted by the model (parameter $\lambda$ in the Poisson model) with the number of days estimated by the expert for a small set of test cases which were not seen during training. The results are shown in Table 9 in Appendix A.
@@
Sampling  The joint probability distribution from Equation (1) is now fully specified. We can use this Bayesian network to randomly sample the tabular portion of a patient record top-down, starting from the root variables without parents at the top and continuing further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process $10{,}000$ times, leaving us with $10{,}000$ artificial patient records consisting of 16 tabular features.
@@ 2.2 Generating unstructured text with a large language model
Once the tabular patient record has been generated, we prompt a large language model (LLM, in our case GPT-4o) to write a clinical note based on the tabular variables associated with this fictional encounter. The scenario we simulate artificially is as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, and other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
@@
To generate such a clinical note, we use a prompt which is generally structured like the example shown in Fig. 2. Additional example prompts can be found in Appendix C. Only the background variables and symptoms may be directly mentioned in the prompt, while the diagnoses pneumonia and common cold would not yet be known to a clinician who is taking notes during a consultation and are therefore left out. Realistically, they can still influence the content of the note through the descriptions of the symptoms that are included in the prompt, as will be explained later. The treatment and outcome are left out of the prompt as well, just like the non-clinical variables, as all of these are typically either unknown or irrelevant at the time of writing the note.We now describe the different parts of the prompt.
Symptom Cause Descriptors
dyspnea asthma attack-related, at night, in episodes, wheezing, difficulty breathing in, feeling of suffocation, nighttime stuffiness, provoked by exercise, light, severe, not able to breathe properly, air hunger
smoking during exercise, worse in morning, mild
COPD chronic, worse during flare-up, worse when lying down, difficulty sleeping, air hunger
hay fever light, mild, stuffy feeling, all closed up
pneumonia light, mild, severe, no clear cause
cough asthma attack-related, dry
smoking productive, mostly in morning, during exercise, gurgling
COPD phlegm, sputum, gurgling, worse when lying down
pneumonia for over 7 days, light, mild, severe, non-productive at first and later purulent
common cold prickly, irritating, dry, phlegm, sputum, light, mild, severe, constant, day and night
pain asthma tension behind sternum
COPD light, mild
cough muscle pain, burning pain in trachea, burning pain in windpipe, scraping pain in trachea, scraping pain in windpipe
pneumonia light, mild, severe, localized on right side, localized on left side, associated with breathing
common cold burning pain in trachea, burning pain in windpipe, scraping pain in trachea, scraping pain in windpipe, light, mild
Special case: no respiratory symptoms  Around one third of all patients in our dataset (3629 out of 10,000 patients sampled from the BN) do not experience any respiratory symptoms at all (all symptoms are “no”). If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example is shown in Fig. 5 in Appendix B.
In these cases, it makes more sense to assume that the patient visits the doctor because of a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Fig. 6 in Appendix B. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, as shown in Fig. 7 in Appendix B. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated notes to each tabular patient record within this subset.
Consistency check  We add an additional filtering step to the generation process to better ensure the quality of the generated notes. We prompt the LLM once again, showing it the same information on the patient’s symptoms and underlying health conditions as in the previous prompt from Fig. 2. We then show the LLM the generated note, and ask it to decide whether it is consistent with this information or not. We ask it to provide an explanation if it deems the note inconsistent. This way, the LLM identified 52 out of $10{,}000$ notes that were inconsistent with the prompt. After manually verifying that this was indeed the case, we asked the LLM to regenerate these notes using the same prompt as before, afterwards finding that none of the newly generated notes were flagged as inconsistent anymore.
@@
Compact version  Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we ask the LLM to create a compact version of each note through an additional prompt, as can be seen in Fig. 2. Our dataset contains both the original note and the compact version of the note.
Prompting details  As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 [21]. We set the temperature to 1.2 to encourage some more variation in the notes. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were set to their default values. Generating all 10,000 notes and their compact version cost around $150$\$.
@@ 2.3 Automated extraction of symptom spans
a LLM-based annotation of a normal clinical note with symptom spans. The model is prompted to return a JSON object of the form {‘‘symptom’’: symptom, ‘‘text’’: extracted phrase}. For clarity, we highlight the extracted spans in the note rather than displaying the raw JSON. Full task instructions are in Appendix D.
b LLM-based annotation of a compact clinical note using spans previously extracted from the corresponding normal note. Full task instructions are provided in Appendix D.
@@
Normal notes Fig. 4a shows the prompt and the extracted phrases for one of the normal notes. We first tell the LLM what symptoms the patient does and does not have, and then ask it to extract phrases that describe these symptoms. We use the system message to send extensive task instructions to the LLM, which we have included in Appendix D. We further process the extracted phrases by matching them to the note, disregarding capital letters and punctuation. If no exact match is found, we use another LLM call to correct the mistake, asking it to map the extracted phrase to an exact phrase in the text.
@@
Dyspnea Cough Pain Nasal Fever All
@@
% of notes where sympt. mentioned ($\geq$ 1 span / note) Normal Compact 99.1% 98.1% 66.5% 62.5% 97.0% 95.4% 64.5% 59.4% 81.7% 79.4 % 39.9% 37.1% 95.7% 95.4% 26.6% 25.1% 95.9% 95.7% 71.6% 71.4% 44.4% 42.4%
Avg. # spans where sympt. mentioned / note Normal Compact 2.03 1.91 0.70 0.65 1.78 1.66 0.65 0.60 1.28 1.21 0.40 0.37 2.47 2.38 0.30 0.28 2.46 2.33 1.55 1.48 0.52 0.48
3 Utility and discussion
We now present a series of experiments to illustrate the utility and potential of our dataset. First, Section 3.1 reports the results of an expert evaluation study assessing the quality of the artificial clinical notes generated by the LLM. Next, Section 3.2 evaluates the performance of simple symptom prediction models applied to both the tabular and textual portions of the dataset. Finally, Section 3.3 explores the broader research applications of the dataset, including but not limited to clinical information extraction.
3.1 Expert evaluation
We asked five general practitioners to evaluate the quality of a random sample of $30$ generated notes. The evaluators got to see the two versions of the generated note (normal and compact), as well as the prompt that was used to generate them. They were asked to evaluate four aspects:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
To measure consistency, the evaluators were asked to assign a penalty for every element in the note that does not correspond with the requested information in the prompt, and for every requested symptom that was missing from the generated note. These penalties were then turned into ratings from 1 (>3 penalties) to 5 (0 penalties). All other aspects were directly rated on a scale of 1 (very bad) to 5 (perfect). Table 4 shows the results. For each rated aspect, we calculate the average score over all $30$ notes, and report the mean and standard deviation over the five evaluators. For more details on the meaning of each rating in each aspect, as well as inter-annotator agreement scores, we refer to Appendix E.
Normal Compact
Consis- tency Realism (hist) Realism (phys) Clinical accuracy Content Reada- bility
mean 4.69 4.53 4.15 4.92 4.88 4.02
std 0.12 0.21 0.30 0.07 0.10 0.31
Results  Our artificial notes were rated as highly consistent with the prompt, and therefore with the information present in the tabular portion of the dataset. As shown in Appendix E.2, a large majority of the inconsistencies arise due to a violation of the additional instructions (usually by inventing additional symptoms), while the key information included in the prompt was still conveyed correctly in the note. High consistency between the tabular variables and the concepts mentioned in the text makes our dataset a reliable resource for information extraction tasks.
The evaluators also deem the notes sufficiently realistic, though the realism of the history section is rated higher on average than the realism of the physical examination. Out of those notes that scored worse, many included a clinical test that seemed unnecessary to the evaluators, while a few forgot a test that was deemed important. At the same time, the very high score for clinical accuracy is an important indication that the notes do not contain falsehoods.
Multiple evaluators mentioned that while the content of the notes seemed realistic, the format did not, as their own notes would be more complex as opposed to the artificial notes, which use clean language and full sentences. This underscores the fact that the dataset should not be used to train any systems which will later be deployed on real notes, and should instead fulfill the role of a research benchmark only.
@@ 3.2 Symptom predictor baselines
In order to set a baseline for future information extraction tasks, we run various prediction models on both the tabular and textual parts of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal. Two of our baselines only get to see the tabular portion of the dataset at the input: Bayesian network (BN-tab) and XGBoost (XGBoost-tab). We use these models to predict each symptom in three settings, differing from one another in the set of tabular features that are taken as an input, which we call the evidence:
@@
[…]
  • $\mathcal{P}(\text{{sympt}}\mid\text{{realistic}})$: Predict the symptom given a more realistic set of tabular features as evidence. We do not expect policy, self-employed and #days to be recorded in any kind of realistic patient record, and therefore leave them out of this evidence set. As in the no-sympt setting, we do not include the symptoms either. In other words, this set includes the background, diagnoses, season and treatment variables.
Apart from the tabular-only baselines, we also train baselines that get to see the text. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. We extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between the tabular concepts, as envisioned in Fig. 1.
Section 3.2.1 provides a detailed overview of each baseline model, followed by Section 3.2.2, which analyzes their performance on the task of automated symptom extraction using the SimSUM dataset.
@@ 3.2.1 Models
BN-tab We provide the causal structure in Fig. 2 to the BN, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters. Where we don’t directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and #days), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix F.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned BN to predict each of the symptoms, taking different sets of variables as evidence according to the three settings described earlier (all, no-sympt and realistic).
XGBoost-tab We train an XGBoost classifier for each symptom in combination with each of the three evidence settings, meaning each classifier sees a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix F.2.
Neural-text We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD-2023 [26]. We explore four settings for turning these sentence embeddings into a single note embedding:
@@
Additionally, we create symptom-specific span embeddings as well. To this end, we take the set of spans mentioning the symptom (as extracted automatically in Section 2.3), transform each of these into their BioLORD embedding, and take the average to get one span embedding per symptom.
No matter the type, each note embedding is fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom $fever$. We trained each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix F.3.
Neural-text-tab  We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed to a one-hot encoding, while the variable #days was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom combined with each evidence setting (all, no-sympt and realistic). For more details, we refer to Appendix F.4.
Setup  We use a random $8{,}000$/$2{,}000$ split to obtain a train and test subset of SimSUM. We use cross-validation on the train set to tune any hyperparameters associated with each model, and report the average F1-score over the test set after training with 5 different seeds [^2] Seeding randomizes the initialization of the model weights, as well as batching.. For the binary symptoms, we use a $0.5$ decision threshold. For fever, which has three possible categories, the class with the highest predicted probability is chosen.
3.2.2 Results

### v4

@@ 2.1 Modeling structured tabular variables with a Bayesian network
The coefficients for each model were tuned using gradient descent based on the train cases shown in Table 7 in Appendix A. The expert was asked to (loosely) label these cases with how long they suspected the patient to stay home as a result of these symptoms. The coefficient for the variable self-employed was tuned manually, based on the assumption that being self-employed would shave some days off the predicted number, regardless of the particular symptoms experienced by the patient. As a sanity check, we compared the mean number of days predicted by the model (parameter $\lambda$ in the Poisson model) with the number of days estimated by the expert for a small set of test cases which were not seen during training. The results are shown in Table 8 in Appendix A.
@@
Sampling  Once the joint probability distribution from Equation (1) is fully specified, we can use this Bayesian network to randomly sample the tabular portion of a patient record top-down. Hereby, we start from the root variables without parents at the top and continue further down. Each value is sampled conditionally on the variable’s parents’ values, using the conditional distributions we have defined. We repeat this process 10,00010{,}000 times, leaving us with 10,00010{,}000 artificial patient records consisting of 16 tabular features.
@@ 2.2 Generating unstructured text with a large language model
Once the tabular patient record has been generated, we prompt a large language model (LLM, in this case GPT-4o) to write a clinical note based on this fictional patient encounter and its associated tabular variables. Only the background variables and symptoms may be directly mentioned in the prompt, while the diagnoses pneumonia and common cold would not yet be known to a clinician who is taking notes during a consultation and are therefore left out. Realistically, they can still influence the content of the note through the descriptions of the symptoms that are included in the prompt, as will be explained later. The treatment and outcome are left out of the prompt as well, just like the non-clinical variables, as all of these are typically either unknown or irrelevant at the time of writing the note.
The prompt is generally structured like the example shown in Fig. 3, and is made up of the following parts. Further details and additional example prompts can be found in Appendix B.1.
  • First, we list the symptoms which are experienced by the patient. We do not exhaustively list the full set of symptoms, but rather sample the probability that a symptom is mentioned according to a distribution defined by the expert. This renders the notes more realistic, since real notes do not exhaustively mention the presence or absence of all possible symptoms, instead following the narrative of the patient and the subsequent probing of the clinician.
  • While we encourage the LLM to invent additional context around the patient’s symptoms, we want this context to at least partially relate to the underlying cause of the symptom. For this reason, we include a descriptor next to each symptom, which is randomly sampled from a set of expert-defined phrases that describe the symptom in the case where it results from a particular cause. This way, the diagnoses can influence the content of the note, even when they are not mentioned explicitly in the prompt. The full list of descriptors, as well as further information on how we sample them, is included in Table 9.
  • The next part of the prompt lists the underlying health conditions, which are assumed to be known up-front, and could therefore influence the content of the note.
  • At the end, the prompt lists some additional instructions. We tell the LLM that the note must be structured with a “History” portion and a “Physical examination” portion. We also ask it not to mention any suspicions of possible diagnoses, nor patient gender or age. Further motivation for these requests can be found in Appendix B.1.
Furthermore, we introduce two additional prompts to extend our dataset and make the notes more realistic.
@@
  • Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we ask the LLM to create a compact version of each note through an additional prompt, as can be seen in the bottom part of Fig. 3.
  • We devise an alternative prompting strategy which is used to generate clinical notes in the special case where no respiratory symptoms are present in the patient record. This occurs for around one third of our dataset. In these cases, we encourage the LLM to imagine an alternative reason for the patient visit, unrelated to the respiratory domain. Examples and further details relating to this alternative strategy are included in Appendix B.3.
According to the prompts outlined above, we ask the LLM to generate both a clinical note and its compact version for each of the 10,00010{,}000 patients in the dataset. After generating each note, we use the LLM to perform an additional consistency check to ensure the quality of the generated notes. More specifically, we prompt the LLM once again, showing it the same information on the patient’s symptoms and underlying health conditions as in the original prompt. We then ask the LLM to decide whether the generated note is consistent with this information or not. This way, the LLM identified 52 out of 10,00010{,}000 notes as inconsistent with the prompt, which were regenerated and manually checked.
@@ 2.3 Automated extraction of symptom spans
a LLM-based annotation of a normal clinical note with symptom spans. The model is prompted to return a JSON object of the form {‘‘symptom’’: symptom, ‘‘text’’: extracted phrase}. For clarity, we highlight the extracted spans in the note rather than displaying the raw JSON. Full task instructions are in Appendix C.
b LLM-based annotation of a compact clinical note using spans previously extracted from the corresponding normal note. Full task instructions are provided in Appendix C.
@@
Normal notes Fig. 5a shows the prompt and the extracted phrases for one of the normal notes. We first tell the LLM what symptoms the patient does and does not have, and then ask it to extract phrases that describe these symptoms. We use the system message to send extensive task instructions to the LLM, which we have included in Appendix C. We further process the extracted phrases by matching them to the note, disregarding capital letters and punctuation. If no exact match is found, we use another LLM call to correct the mistake, asking it to map the extracted phrase to an exact phrase in the text.
@@
Using the strategy outlined above, we extract all text spans mentioning each of the 5 symptoms from all 10,00010{,}000 notes in our dataset. In Section 3.2, we use these fine-grained extracted spans to analyze the content of the notes and quantify the presence of each symptom in the textual portion of SimSUM.
3 Utility and discussion
We now present a series of experiments to illustrate the utility and potential of our dataset. First, Section 3.1 reports the results of an expert evaluation study assessing the quality of the artificial clinical notes generated by the LLM. Next, Section 3.2 uses the extracted symptom spans to further analyze the content of the notes. Following this, Section 3.3 evaluates the performance of simple symptom prediction models applied to both the tabular and textual portions of the dataset. Finally, Section 3.4 explores the broader research applications of the dataset, including but not limited to clinical information extraction.
3.1 Expert evaluation of generated notes
We asked five general practitioners to evaluate the quality of a random sample of $30$ generated notes. The primary aim of our evaluation study was to assess whether the generated notes adhered to the instructions provided in the prompt. As a secondary aim, we sought to gain an indication of how realistic the content of the notes appeared. We invited general practitioners with an affiliation to the Department of Primary Care at our institution to take part in the evaluation study. The participants were recruited via email, and included based on their willingness to participate when provided information about the task’s duration. The evaluators got to see the two versions of the generated note (normal and compact), as well as the prompt that was used to generate them.
3.1.1 Evaluation criteria
The evaluators were asked to rate four criteria for each clinical note:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
To measure consistency, the evaluators were asked to assign a penalty for every element in the note that does not correspond with the requested information in the prompt, and for every symptom requested in the prompt that was missing from the generated note. These penalties were then turned into ratings from 1 (>3 penalties) to 5 (0 penalties). All other aspects were directly rated on a scale of 1 (very bad) to 5 (perfect). To ensure the clarity and intuitiveness of the evaluation criteria, we conducted a pilot run of the evaluation process and obtained feedback from a general practitioner who did not participate in the final study. For more details on the meaning of each rating in each criterion, as well as inter-annotator agreement scores, we refer to Appendix D.
3.1.2 Evaluation results
Normal Compact
Consis-
tency
Realism
Realism
Clinical
accuracy
Content
Reada-
bility
mean 4.69 4.53 4.15 4.92 4.88 4.02
std 0.12 0.21 0.30 0.07 0.10 0.31
Table 2 shows the results of the evaluation study. For each rated criterion, we calculate the average score over all $30$ notes, and report the mean and standard deviation over the five evaluators. Our artificial notes were rated as highly consistent with the prompt, and therefore with the information present in the tabular portion of the dataset. As shown in Appendix D.2, a large majority of the inconsistencies arise due to a violation of the additional instructions (usually by inventing additional symptoms), while the key information included in the prompt was still conveyed correctly in the note. High consistency between the tabular variables and the concepts mentioned in the text makes our dataset a reliable resource for information extraction tasks.
The evaluators also deem the notes sufficiently realistic, though the realism of the history section is rated higher on average than the realism of the physical examination. Out of those notes that scored worse, many included a clinical test that seemed unnecessary to the evaluators, while a few forgot a test that was deemed important. At the same time, the very high score for clinical accuracy is an important indication that the notes do not contain falsehoods. Multiple evaluators mentioned that while the content of the notes seemed realistic, the format did not, as their own notes would be more complex as opposed to the artificial notes, which use clean language and full sentences. This underscores the fact that the dataset should not be used to train any systems which will later be deployed on real notes, and should instead fulfill the role of a research benchmark only.
3.2 Analysis of generated notes with symptom spans
@@
Dyspnea
Cough
Pain
Nasal
Fever All
@@
% of notes where
sympt. mentioned
($\geq$ 1 span / note)
Normal
Compact
99.1%
98.1%
66.5%
62.5%
97.0%
95.4%
64.5%
59.4%
81.7%
79.4 %
39.9%
37.1%
95.7%
95.4%
26.6%
25.1%
95.9%
95.7%
71.6%
71.4%
44.4%
42.4%
Avg. # spans where
sympt. mentioned / note
Normal
Compact
2.03
1.91
0.70
0.65
1.78
1.66
0.65
0.60
1.28
1.21
0.40
0.37
2.47
2.38
0.30
0.28
2.46
2.33
1.55
1.48
0.52
0.48
The symptom spans extracted in Section 2.3 allow us to quantify the presence of each of the 5 symptoms in the notes. We first evaluate the accuracy of the extracted spans with a small manual evaluation, after which we use the spans to analyze the content of the notes.
@@ 3.3 Symptom predictor baselines
In order to set a baseline for future information extraction tasks, such as the one envisioned in Fig. 2, we run various prediction models on both the tabular and textual parts of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal. Symptom prediction is a useful task to show the utility of our dataset, since information about the symptoms can be inferred from both the tabular portion of the dataset (through the other variables), as well as the textual portion of our dataset (the clinical notes which partially describe the symptoms).
Section 3.3.1 describes the task setup and gives an overview of the various types of baselines we run on our dataset. Section 3.3.2 provides a detailed overview of each baseline model, followed by Section 3.3.3, which analyzes their performance on the task of automated symptom extraction using the SimSUM dataset.
3.3.1 Task overview
Tabular-only baselines  Two of our baselines only get to see the tabular portion of the dataset at the input: Bayesian network (BN-tab) and XGBoost (XGBoost-tab). We include XGBoost because it has consistently demonstrated state-of-the-art performance on a wide range of tabular prediction tasks, often serving as a strong baseline in machine learning benchmarks [xgboost_tabular]. We use both models to predict each symptom in three settings, differing from one another in the set of tabular features that are taken as an input, which we call the evidence:
@@ 3.3.1 Task overview
  • $\mathcal{P}(\text{{sympt}}\mid\text{{no-sympt}})$: Predict the symptom given all other tabular features as evidence, except for the other symptoms. This mimics the setting where we have tabular features available in the patient record, but have not extracted any symptoms from the text yet. This set includes the background, diagnoses, non-clinical, treatment and outcome variables.
  • $\mathcal{P}(\text{{sympt}}\mid\text{{realistic}})$: Predict the symptom given a more realistic set of tabular features as evidence. We do not expect policy, self-employed and days at home to be recorded in any kind of realistic patient record, and therefore leave them out of this evidence set. As in the no-sympt setting, we do not include the symptoms either. In other words, this set includes the background, diagnoses, season and treatment variables.
Text baselines  Apart from the tabular-only baselines, we also train baselines that get to see the text. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. We extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between the tabular concepts, as envisioned in Fig. 2.
@@ 3.3.2 Models
BN-tab We provide the causal structure in Fig. 3 to the BN, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters. Where we do not directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and days at home), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix E.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned BN to predict each of the symptoms, taking different sets of variables as evidence according to the three settings described earlier (all, no-sympt and realistic).
XGBoost-tab We train an XGBoost classifier for each symptom in combination with each of the three evidence settings, meaning each classifier sees a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix E.2.
Neural-text We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD-2023 [biolord]. We explore four settings for turning these sentence embeddings into a single note embedding:
@@
Additionally, we create symptom-specific span embeddings for each note as well. To this end, we take the set of spans mentioning the symptom in each note (as extracted automatically in Section 2.3), transform each of these spans into their BioLORD embedding, and take the average to get one span embedding per symptom per note. No matter the type, each note embedding is fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom feverfever. We trained each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix E.3.
Neural-text-tab  We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed to a one-hot encoding, while the variable days at home was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom combined with each evidence setting (all, no-sympt and realistic). For more details, we refer to Appendix E.4.
Training setup  We use a random 8,0008{,}000/2,0002{,}000 split to obtain a train and test subset of SimSUM. We use cross-validation on the train set to tune any hyperparameters associated with each model, and report the average F1-score over the test set after training with 5 different seeds. Seeding randomizes the initialization of the model weights, as well as batching. For the binary symptoms, we use a $0.5$ decision threshold. For fever, which has three possible categories, the class with the highest predicted probability is chosen.
3.3.3 Symptom predictor results